OpenStudy (anonymous):

factor completley 2r-6-6r+18

OpenStudy (anonymous):

Answer choices???

OpenStudy (anonymous):

solve by grouping...based upon your other questions. so let's rewrite to regroup variables that are related.

OpenStudy (anonymous):

2r-6-6r+18 = 0 2r -6r -6 +18 = 0 -4r +12 = 0 -4r = -12 r = 4 actually you could just combine like terms. Don't forget the =0 that is very important because with out you are not able to "solve"

OpenStudy (anonymous):

Or if you wrote the problem wrong...with "factoring" I think you did..check that please

OpenStudy (anonymous):

2r-6-6r+18 = 2r-6r-6+18 = 2r(1-3)-6(1-3) = (2r-6)(1-3)

OpenStudy (anonymous):

agreed!

OpenStudy (anonymous):

you can re-check the factor (2r-6)(1-3)... you'll get the same